Circular No. 7843

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SAX J1805.5–2031 P. Lowes, National Institute for Space Research, Space Research Orga-

nization Netherlands (SRON), Utrecht, and BeppoSAX Science Operation Center (BSOC), Rome; J. J. M. in ’t Zand, Astronomical Institute, Utrecht University, and SRON; J. Heise, SRON; M. Cocchi and L. Natalucci, Istituto di Astrofisica Spaziale e Fisica Cosmica, Rome; G. Gennaro and M. Stornelli, BSOC, report that on Mar. 5.5–5.9 UT, Wide Field Camera unit 1 on BeppoSAX detected a new x-ray transient, designated SAX J1805.5–2031 and located at α = 18^{h}05^{m}34^{s}, δ = −20^{o}30′.8 (equinox 2000.0; uncertainty 1′.8, 99-percent confidence). The flux was 0.04 Crab units in the range 2–28 keV. No x-ray bursts were detected.

C. B. Markwardt, University of Maryland and Goddard Space Flight Center (GSFC); and E. Smith and J. Swank, GSFC, report that the source SAX J1805.5–2031 was observed by the RXTE PCA at 2–10 keV in monitoring scans of the Galactic bulge region around Mar. 2.97 UT (flux 51 ± 2 mCrab) and in a 3100-s pointed PCA observation starting Mar. 5.84 (flux 36 ± 1 mCrab). The position of the source in the PCA scans is consistent with the BeppoSAX WFC position (cf. Lowes et al., above). The 2–20-keV spectrum is well fitted with a combination of blackbody (kT = 0.9 keV) emission and a power law (photon index −2.9) for a column density of 4 × 10^{22} cm^{−2}. No strong pulsations of quasiperiodic oscillations are detected; the total fractional r.m.s. variability of the source is 16 percent in the band 0.001–2048 Hz. Additional RXTE observations are planned.

COMET C/2002 C1 (IKEYA-ZHANG) B. G. Marsden (cf. MPEC 2002-D36) and S. Nakano have noted the

similarity of the orbits of comets C/2002 C1 and C/1661 C1 and that numerical integration of the 2002 orbit backward yields a previous perihelion date within a couple of years of 1659, making the link rather likely. Additional selected visual m_{1} estimates: Feb. 14.74 UT, 7.1 (H. Dahle, Copenhagen, Denmark, 9×63 binoculars); 18.95, 6.8 (J. G. de S. Aguiar, Campinas, Brazil, 11×80 binoculars); 20.10, 6.4 (A. Hale, Cloudcroft, NM, 11×80 binoculars); 21.77, 6.3 (A. Kammerer, Ettlingen, Germany, 9×63 binoculars); 24.76, 6.0 (M. V. Zanotta, Laino, Italy, 7×42 binoculars); 28.78, 5.3 (R. J. Bouma, Groningen, The Netherlands, 15×80 binoculars); Mar. 3.75, 4.8 (K. Hornoch, Lelekovice, Czech Republic, naked eye); 4.41, 5.2 (Y. Nagai, Yamanashi, Japan, 12×50 binoculars); 6.49, 5.2 (A. Pearce, Nedlands, W. Australia, 20×80 binoculars).

2002 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green